$SPAD/src/input f04faf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)cle all
showArrayValues true
showScalarValues true
job:=0
n:=5
d:Matrix SF:=
 [[4 ,10 ,29 ,25 ,5 ]]
e:Matrix SF:=
 [[0,-2 ,-6 ,15 ,8 ]]
b:Matrix SF:=
 [[6 ,9 ,2 ,14 ,7 ]]
result:=f04faf(job,n,d,e,b,-1)

———-

References^{[1] nothing}

4